# Synthesizing Efficient and Permissive Programmatic Runtime Shields for Neural Policies

### v1

@@ 5. Discussion
The figure shows that neural policies consistently require fewer steps to reach steady states compared to our runtime shields, with an average of 10.5 steps for neural policies versus 19.3 steps for runtime shields. This suggests that neural policies are more efficient at maintaining system stability. Therefore, we conclude that neural policies are essential for control systems, while runtime shields serve primarily to ensure the safety of these policies. Additionally, we observe that neural policies paired with Aegis require more steps to reach steady states than those without shields, increasing the average steps from 10.5 to 15.75. This demonstrates that deploying runtime shields can compromise system stability, reflecting the “no free lunch” theorem (Wolpert and Macready, 1997), where safety comes at the cost of stability. This observation somewhat aligns with the finding of Zhu et al. (Zhu et al., 2019), which highlights that while alternative synthesis methods such as LQR-based approaches (Tedrake, 2010; Reist et al., 2016) can yield controllers with superior stability, they often fail to meet safety constraints. Our approach, therefore, represents a trade-off between safety and stability, but prioritizing safety. We leave the exploration of methods that can achieve better stability while maintaining safety to future work.
Figure 8. Time steps to reach steady state for the quadcopter system.
@@ 6. Related Work
Runtime shields have been used in runtime enforcement for reactive systems (Wu et al., 2019; Könighofer et al., 2017; Akametalu et al., 2014; Gillula and Tomlin, 2012; Bloem et al., 2015; Könighofer et al., 2020; Alshiekh et al., 2018) long before the advent of neural policies. However, runtime shields for neural policies are a relatively new concept and have only been studied in a few papers. Runtime shields in reinforcement learning models were first introduced in (Alshiekh et al., 2018). Because their verification approach is not symbolic, however, it can only work over finite discrete states instead of continuous state settings like the ones we consider in this paper. Zhu et al. (Zhu et al., 2019) proposed a method to synthesize deterministic and verified programs from neural policies, which can be used as runtime shields for neural policies. However, their shields have limitations in terms of efficiency and permissiveness, which necessitates the need for a more effective and permissive approach, motivating our work. Xiong et al. (Xiong and Jagannathan, 2021) proposed a method capable of synthesizing high-quality runtime shields even when the subject system involves hundreds of dimensions. However, their method is not effective enough to synthesize runtime shields that can guarantee safety for neural policies but still occasionally let unsafe commands pass through. Therefore, we do not consider it as a baseline in our experiments. Our approach differs from these studies in that we propose a new approach that can synthesize shields with lightweight overhead and high permissiveness, which is a novel and unique perspective in the context of runtime shields for neural policies.
@@
Although program synthesis has been studied for decades (Sobania et al., 2022), very little attention has been paid to program synthesis specialized for neural policies. A few studies (Verma et al., 2018; Trivedi et al., 2021) have proposed methods for generating programmatic policies from well-trained DRL models, which are more interpretable than neural networks. Additionally, Inala et al. (Inala et al., 2020) and Qiu et al. (Qiu and Zhu, 2021) synthesize programmatic controllers to perform continuous control tasks. Moreover, there are a few studies that focus on synthesizing code implementations or configurations for deep learning models in various tasks (Zhou et al., 2024; Shi et al., 2024; Liu et al., 2023; Shi et al., 2023). Our approach differs from these studies in that we propose a counterexample-guided inductive program synthesis approach which can efficiently synthesize runtime shields for neural policies, which is a unique application of program synthesis for neural policies.
@@
Formal verification of neural policies has been explored in several studies (Landers and Doryab, 2023; Mazouni et al., 2023). Some of these studies align with our work by using counterexample-guided inductive synthesis (CEGIS) to formally verify neural policies and their neural certificates. Certain works, including (Chang et al., 2019; Abate et al., 2020, 2021; Edwards et al., 2024), focus on deterministic systems, employing symbolic reasoning to verify neural policies and construct Lyapunov functions (Haddad and Chellaboina, 2008) or barrier certificates in a CEGIS loop. Another more recent line of work (Lechner et al., 2022; Žikelić et al., 2023; Chatterjee et al., 2023; Ansaripour et al., 2023) considers both deterministic and stochastic systems, utilizing discretization and abstract interpretation for verification within the CEGIS loop, and deriving neural certificates of correctness or stability in the form of supermartingales instead of Lyapunov and barrier functions. Our work, however, differs by focusing on synthesizing runtime shields as simple programs for neural policies. This approach aims to enhance the efficiency and permissiveness of these shields, leading to a unique application of formal verification in the context of neural policies.
@@ 7. Conclusion and Future Work
In this paper, we present Aegis, a novel approach to synthesizing lightweight and permissive programmatic runtime shields for neural policies. Aegis achieves this by formulating the seeking of a runtime shield as a sketch-based program synthesis problem and using two novel algorithms to solve it, namely a counterexample-guided inductive synthesis algorithm and a Bayesian optimization algorithm. To evaluate Aegis and its synthesized shields, we use four representative control systems with neural policies and compare Aegis to the current state-of-the-art VRL. Our results show that the programmatic runtime shields synthesized by Aegis can correct all unsafe commands from neural policies, ensuring that the systems do not violate the desired safety properties at any time. In particular, compared to VRL, Aegis’s runtime shields exhibit a 2.1$\times$ reduction in time overhead and a 4.4$\times$ reduction in memory overhead, suggesting that they are much more lightweight. Additionally, Aegis’s shields show improved permissiveness, leading to an average of 1.6$\times$ fewer interventions compared to VRL’s shields. These results not only highlight the superiority of Aegis over the state-of-the-art, but also show that Aegis is an effective framework for synthesizing programmatic runtime shields for neural policies.
In the future, we plan to extend Aegis to handle more complex control systems such as those that take more complex inputs, including images and videos. We also plan to work on more advanced program synthesis techniques to further improve Aegis.
Replication Package.   For reproducibility and advancing the state of research, we have made our tool and all experimental data publicly available: https://github.com/jiekeshi/AEGIS.

### v3

@@ 5. Discussion
QuadcopterPendulum-v1Pendulum-v2Cart-Pole-v1Cart-Pole-v2Self-Driving-v1Self-Driving-v24-Car Platoon01530456022437179143024158252118315928871619183360Neural Policy (avg. 13.3) Aegis (avg. 25.1) Neural Policy w/ Aegis (avg. 23.6)
Figure 14. Time steps to reach steady states.
@@
The figure shows that neural policies consistently require fewer steps to reach steady states compared to our runtime shields, with an average of 13.3 steps for neural policies versus 25.1 steps for runtime shields. This suggests that neural policies are more efficient at maintaining system stability. Therefore, we conclude that neural policies are essential for control systems, while runtime shields serve primarily to ensure the safety of these policies. Additionally, we observe that neural policies paired with Aegis require more steps to reach steady states than those without shields, increasing the average steps from 13.3 to 23.6. This demonstrates that deploying runtime shields can compromise system stability, reflecting the “no free lunch” theorem (Wolpert and Macready, 1997), where safety comes at the cost of stability. This observation somewhat aligns with the finding of Zhu et al. (Zhu et al., 2019), which highlights that while alternative synthesis methods such as LQR-based approaches (Tedrake, 2009; Reist et al., 2016) can yield controllers with superior stability, they often fail to meet safety constraints. Our approach, therefore, represents a trade-off between safety and stability, but prioritizing safety. We leave the exploration of methods that can achieve better stability while maintaining safety to future work.
5.2. Ablation Study
Table 6. Ablation study of Aegis.
# Safety Violations Time Overhead Memory Overhead # Shield Interventions
Aegis 0/1000 (100% Safe) 1.28s 0.85 MB 313
w/o Synthesis 43/1000 (4.3% Unsafe) 2.10s 0.88 MB 390
w/o Optimization 117/1000 (11.7% Unsafe) 3.99s 0.85 MB 642
We perform an ablation study by comparing the performance of Aegis with two variants, each lacking a key component. Table 6 reports the average results across all eight control systems on the main evaluation metrics. Specifically, these two variants are:
w/o Synthesis
This variant omits the counterexample-guided inductive synthesis process (Section 3) and instead directly uses a Linear-Quadratic Regulator (LQR) to initialize a linear policy. This policy is then paired with the optimized switching strategy to construct the runtime shield. However, without counterexample-driven refinement, the policy is not guaranteed to be safe across the full state space, as it is not iteratively improved using concrete counterexamples. As shown in Table 6, this variant results in an average of 43 safety violations—substantially more than the 0 violations achieved by the full version of Aegis. Although time and memory overheads and intervention counts are only moderately affected, this is expected: time overhead and interventions are primarily governed by the switching strategy, and memory usage is dominated by the linear policy structure. These findings underscore the importance of the synthesis process for ensuring robust safety guarantees.
w/o Optimization
[…]
5.3. Fidelity of Inferred Environment Dynamics
Our synthesis procedure assumes that the environment dynamics can be accurately inferred. To assess this assumption, we conduct experiments comparing the outputs of the inferred dynamics with those of the actual ones. Following a similar evaluation setup as in our RQs, we run each control system for 5,000 time steps using both the original and inferred dynamics. We collect the resulting state trajectories and quantify their difference using the mean squared error (MSE), defined as:
(30) $\text{MSE}=\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}$
where $y_{i}$ is the actual state at time step $i$, $\hat{y}_{i}$ is the predicted state from the inferred dynamics, and n=5,000n=5{,}000 is the number of samples. A lower MSE indicates higher fidelity between the inferred and actual dynamics. The results are presented in the first column of Table 7. In our experiments, we use the initial state of the system, which is assumed to be within a neighborhood of the equilibrium point, as the input for inferring environment dynamics. The results show an average MSE of only 3.2e-8, with a maximum MSE of 9.9e-8, both of which are extremely low. These small error values suggest that the inferred dynamics closely approximate the true system dynamics, indicating that any adverse effects on the shield synthesis process are likely negligible.
Table 7. The mean absolute error (MAE) between the inferred and actual environment dynamics.
Control System MAE
random initial state equilibrium point
Quadcopter 1.4e-7 6.2e-10
Pendulum-v1 3.9e-10 1.1e-13
Pendulum-v2 5.6e-9 0.0
CartPole-v1 4.3e-10 0.0
CartPole-v2 2.2e-11 2.3e-11
Self-Driving-v1 2.5e-9 1.8e-9
Self-Driving-v2 6.3e-9 1.8e-9
4-Car Platoon 9.9e-8 0.0
Average 3.2e-8 5.3e-10
Impact of Equilibrium Points
In our experiments, for each run, we randomly sample the initial state of the system over the set of initial states. This approach ensures that the evaluation is not biased toward any specific starting condition. However, as described in Section 3.3, environment dynamics inference tends to be most accurate at equilibrium points where the system is stable and behaves in a locally linear and time-invariant manner. To assess the impact of using equilibrium points, we conduct a supplementary experiment in which we manually compute the equilibrium point for each control system and use it as the initial state for dynamics inference. The results, shown in the second column of Table 7, indicate a further reduction in MSE, with an average value of just 5.3e-10. While this confirms that inference at equilibrium points yields even higher fidelity, the improvement is minimal in absolute terms, with differences appearing only at the eighth decimal place. This suggests that, in our setting, the shield synthesis procedure remains robust even when dynamics are inferred from arbitrary (non-equilibrium) initial states. Since equilibrium points may not be known in advance in practical scenarios, we believe that using randomly sampled initial states for inferring environment dynamics remains a reasonable and effective choice.
5.4. Limitations of Aegis
While the above evaluation demonstrates the effectiveness of Aegis in synthesizing runtime shields for neural policies across diverse control systems, we acknowledge a key limitation: Aegis assumes that safety specifications define convex regions, typically expressed as conjunctions of linear inequalities. This limitation stems from the mathematical foundation of Aegis, which relies on linear programming to compute the maximal output admissible set (see Section 3.4). Disjunctive specifications, which define disconnected or hole-containing regions in the state space, are non-convex and thus fall outside the scope of linear programming-based safety verification. Importantly, this convexity assumption is not unique to Aegis. Prior methods such as VRL (Zhu et al., 2019) make the same assumption, as their synthesis procedure constructs barrier certificates whose safe regions must be convex to enable tractable verification using convex optimization techniques such as sum-of-squares programming.
Despite this theoretical limitation, we argue that many real-world safety constraints—such as bounds on velocity, acceleration, or joint angles—are naturally convex and can be effectively represented by conjunctions of linear inequalities. Therefore, while extending support to non-convex safety specifications remains a promising direction for future work, the current formulation of Aegis already addresses a wide range of practical and safety-critical scenarios and thus remains valuable.
@@ 6. Related Work
Runtime shields have been used in runtime enforcement for reactive systems (Wu et al., 2019; Könighofer et al., 2017; Akametalu et al., 2014; Gillula and Tomlin, 2012; Bloem et al., 2015; Könighofer et al., 2020; Alshiekh et al., 2018; Yang et al., 2021a) long before the advent of neural policies. However, runtime shields for neural policies are a relatively new concept and have only been studied in a few papers. Runtime shields in reinforcement learning models were first introduced in (Alshiekh et al., 2018). Because their verification approach is not symbolic, however, it can only work over finite discrete states instead of continuous state settings like the ones we consider in this paper. Zhu et al. (Zhu et al., 2019) proposed a method to synthesize deterministic and verified programs from neural policies, which can be used as runtime shields for neural policies. However, their shields have limitations in terms of efficiency and permissiveness, which necessitates the need for a more effective and permissive approach, motivating our work. Xiong et al. (Xiong and Jagannathan, 2022) propose a method capable of synthesizing high-quality runtime shields even when the subject system involves hundreds of dimensions. RMPS (Li and Bastani, 2020) achieves runtime shielding via backup model-predictive control (MPC) controllers that can override unsafe commands from neural policies. However, both methods are not sufficiently effective to synthesize runtime shields that can guarantee safety but still occasionally let unsafe commands pass through. Therefore, we do not consider it as a baseline in our experiments.
Several other works aim to enhance the safety of neural policies during training by incorporating runtime shield feedback into the learning loop. For instance, REVEL (Anderson et al., 2020) uses shield interventions as feedback to iteratively refine the learned policy via imitation learning. Similarly, VEL (Wang and Zhu, 2023) leverages verification feedback to iteratively optimize programmatic controllers with provable safety guarantees. In addition to optimization via runtime shield feedback, other approaches improve policy safety through specialized policy optimization techniques, such as policy gradient (Papini et al., 2022). For example, Yu et al. (Yu et al., 2019) propose a convergent policy gradient algorithm for safety-constrained reinforcement learning, while Honari et al. (Honari et al., 2024) introduce a multi-objective optimization approach that balances reward and safety. Verma et al. (Verma et al., 2019) further combine policy gradient methods with program synthesis to learn safe and verifiable policies, with theoretical convergence guarantees. Moreover, a separate line of work considers formal verification and synthesis of neural policies without relying on runtime shielding. This includes methods that apply star set-based reachability analysis (Tran et al., 2019; Lal and Prabhakar, 2023), leverage Lyapunov functions (Zhang et al., 2025; Chow et al., 2018; Chang et al., 2019a; Abate et al., 2021b), or utilize supermartingales (Ansaripour et al., 2023; Chatterjee et al., 2023; Žikelić et al., 2023; Mathiesen et al., 2023) to provide rigorous safety guarantees.
@@
Although program synthesis has been studied for decades (Sobania et al., 2023), very little attention has been paid to program synthesis specialized for neural policies. A few studies (Verma et al., 2018; Trivedi et al., 2021) have proposed methods for generating programmatic policies from well-trained DRL models, which are more interpretable than neural networks. Additionally, Inala et al. (Inala et al., 2020) and Qiu et al. (Qiu and Zhu, 2022) synthesize programmatic controllers to perform continuous control tasks. Moreover, there are a few studies that focus on synthesizing code implementations or configurations for deep learning models in various tasks (Zhou et al., 2024; Shi et al., 2024; Liu et al., 2023; Shi et al., 2023, 2025a). Our approach differs from these studies in that we propose a counterexample-guided inductive program synthesis approach which can efficiently synthesize runtime shields for neural policies, which is a unique application of program synthesis for neural policies.
Another line of work (Mandal et al., 2023a; Liskowski et al., 2020; Ellis et al., 2018; Gaunt et al., 2017) formulates program synthesis as a continuous optimization problem, rather than selecting discrete grammar components as in many syntax-guided approaches (Balog et al., 2017; Bunel et al., 2018). While our method shares this high-level formulation, i.e., seeking optimal parameters in a continuous space, it differs significantly in both objective and technical design. Concretely, we focus on synthesizing runtime shields guided by formal safety constraints, whereas prior work targets functional correctness from input-output examples (Mandal et al., 2023a; Ellis et al., 2021), natural language specifications (Mandal et al., 2023b; Desai et al., 2016), or even visual inputs such as images (Chandra et al., 2022; Ellis et al., 2018). Meanwhile, our approach technically diverges by using refinement based on the maximal admissible set and applying Bayesian optimization, distinguishing it from prior methods that rely on evolutionary strategies (Mandal et al., 2023a; Liskowski et al., 2020) or neural networks (Ellis et al., 2018, 2021).
@@
Formal verification of neural policies has been explored in several studies (Landers and Doryab, 2023; Mazouni et al., 2023). Some of these studies align with our work by using counterexample-guided inductive synthesis (CEGIS) to formally verify neural policies and their neural certificates. Certain works, including (Chang et al., 2019b; Abate et al., 2021b, a; Edwards et al., 2024), focus on deterministic systems, employing symbolic reasoning to verify neural policies and construct Lyapunov functions (Haddad and Chellaboina, 2008) or barrier certificates in a CEGIS loop. Another more recent line of work (Lechner et al., 2022; Žikelić et al., 2023; Chatterjee et al., 2023; Ansaripour et al., 2023) considers both deterministic and stochastic systems, utilizing discretization and abstract interpretation for verification within the CEGIS loop, and deriving neural certificates of correctness or stability in the form of supermartingales instead of Lyapunov and barrier functions. Our work, however, differs by focusing on synthesizing runtime shields as simple programs. This approach aims to enhance the efficiency and permissiveness of these shields, leading to a unique application of formal verification in the context of neural policies.
Linear Constraints of Neural Networks
There is a rich body of work on designing neural networks whose outputs or activations satisfy a set of linear constraints by construction (Qu and Hu, 2011; Hendriks et al., 2021; Frerix et al., 2020; Pathak et al., 2015; Stoian et al., 2024; Tordesillas et al., 2023). Among these, Frerix et al. (Frerix et al., 2020) enforce hard convex constraints on the activations of deep neural networks through a reparameterization of the feasible set. Hendriks et al. (Hendriks et al., 2021) and Tordesillas et al. (Tordesillas et al., 2023) focus on imposing linear constraints on neural network outputs, even under low-data conditions or with very low inference latency. Stoian et al. (Stoian et al., 2024) introduce a constraint layer for generative models that guarantees user-defined linear constraints on tabular data, ensuring that generated samples remain realistic and domain-compliant. In addition, Giunchiglia et al. (Giunchiglia et al., 2023, 2025) recently propose methods and a dataset for training networks that adhere to safety requirements expressed as linear or logical constraints. While these approaches enforce constraints that may be interpreted as safety properties, they assume the network can be retrained and its internal parameters are accessible. In contrast, as discussed in Section 1, our work addresses a fundamentally different setting when retraining the neural policy is prohibitively expensive, or its internal structure is inaccessible, such as in proprietary third-party systems.
@@ 7. Conclusion and Future Work
[…]
In the future, we plan to extend Aegis to support more complex control systems that process richer inputs, such as images (Yang et al., 2022) and audio (Yang et al., 2025a). We also intend to develop more advanced program synthesis techniques to further improve Aegis, for example by incorporating neural-guided synthesis (Barnaby et al., 2025) or Large Language Models (Hou et al., 2024; Yang et al., 2025b; Shi et al., 2022) to enhance the efficiency of the synthesis process.
Replication Package.   For reproducibility and advancing the state of research, we have made our tool and all experimental data publicly available: https://github.com/soarsmu/AEGIS.